OpenStudy (anonymous):

Analyze the function f(x) = - tan 4x. Include: - Domain and range - Period - Two Vertical Asymptotes Thanks so much for your help

OpenStudy (anonymous):

\[Domain { x \in \mathbb{R} : \neq \pi/8 \pm 2\pi/8}\]\[Range { y \in \mathbb{R} }\] You can get the Period from the Domain and the Vertical asymptotes there too.

OpenStudy (anonymous):

I'm sorry could you please explain to me how you got the answer? I'm really trying to learn this stuff I'm just having a tough time. Thanks for answering!

OpenStudy (anonymous):

Graph the function (as can be seen here: http://www.wolframalpha.com/input/?i=f%28x%29+%3D+-+tan+4x) The domain is everything x can be, which is all real numbers, except for -pi/8, pi/8, 3pi/8 (as those are the ends of the periods). The ends of the periods are also the vertical asymptotes. As for the range, y can be any number because it continues upward infinitely.

OpenStudy (anonymous):

So the domain and range the same in this case?

OpenStudy (anonymous):

No, the Domain and Range are different because the domain has asymptotes whilst the range, as Ash has said, can be anything

OpenStudy (anonymous):

So there is no way to narrow down the range in this case? You can say the range is all real numbers?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

no probs =D

OpenStudy (anonymous):

Yeah thanks. So the vertical asymptotes are +- pi/8 and 3pi/8?

OpenStudy (anonymous):

Hmm, just -pi/8 and 3pi/8.

OpenStudy (anonymous):

+/- 2pi/8 as well

OpenStudy (anonymous):

where is the period and the two vertical asymptotes ?